# Quantum resolution limit of long-baseline imaging using distributed entanglement

### v1

@@ II Our physical set-up
where in a nutshell, $|1_{A_{q}},0_{B}\rangle$ and $|0_{A},1_{B_{q}}\rangle$ are the kets associated with a single photon in mode $q$ at sites $A$ and $B$, respectively, and we will make this notation more precise later.
Figure 2: Schematic showing light originating at object plane, propagating until reaching aperture plane, with $n=2$ hard apertures of size $\delta$, and finally producing an image at image plane.
III Joint quantum state of the collected light
We will do this in three steps. First we write the quantum state for a single time bin for just one telescope. We will then give the generalization to two telescopes, and finally, we will write down the full quantum state for multiple time bins.
III.1 The single-aperture case
We take the probability of a single incoming photon in a given temporal mode to be $\epsilon$ Tsang et al. (2016a). The density operator for a thermal source, weak enough to only consider linear terms after expanding in $\epsilon$, is
@@ III.1 The single-aperture case
where $\rho_{0}=|0\rangle\langle 0|$ denotes vacuum and the one-photon state is a mixture of states of the incoming light from each star, labeled $s\in\{1,2\}$:
@@
In the single aperture case, we can write the state for star $s$ as
@@
where $x$ is the coordinate over the image-plane, $|x\rangle=a^{\dagger}(x)|0\rangle$ is the state of a single photon at position $x$, and $a^{\dagger}(x)$ the bosonic creation operator in the continuum limit which satisfies the usual commutation relation $[a(x),a^{\dagger}(x^{\prime})]=\delta(x-x^{\prime})$ Yuen and Shapiro (1978). Lastly, $\psi(x-x_{s})$ is the point-spread function (PSF) corresponding to a single aperture, with $x_{s}$ being the position of the $s$-th star in the object plane.
Now, consider sorting light collected by our aperture in terms of spatial mode basis functions $\phi_{q}(x)$ with $q=0,1,2,\ldots$. Expanding the state $\left|\psi^{(s)}\right\rangle$ defined in Eq. (4) into such a basis gives
@@
is the ket for a single photon in the modal function $\phi_{q}(x)$ (and no photons in all the other modes), and the coefficients $\Gamma_{q}(x_{s})$ are defined as the correlation functions between the PSF and the modal functions:
@@
The probability of measuring an incoming photon in the above basis and finding it to be in the $q$th mode is then given as $p_{q}(x_{s})=|\Gamma_{q}(x_{s})|^{2}$.
To obtain the PSF, we need to take the Fourier transform of the aperture function $\widetilde{\psi}(y)$, where $y$ is the coordinate in the aperture plane. We will consider the realistic scenario of a hard aperture function
@@
where $\delta$ is the size of the aperture. The Fourier transform of this yields the PSF
$\psi(x)=\sqrt{\sigma}\,\frac{\sin(\pi x/\sigma)}{\pi x}.$ (10)
@@
While we can measure in any basis, one natural basis set can be constructed by taking the PSF as the fundamental mode, then taking derivatives of it of increasing order, and finally applying a Gram-Schmidt process to obtain the subsequent modes. For the above PSF for a hard aperture in 1 dimension, this results in the Sinc-Bessel (SB) basis, for which the mode functions are $\phi_{q}(x)\equiv\sqrt{(1+2q)/\sigma}j_{q}(\pi x/\sigma)$, where $j_{q}(\pi x/\sigma)$ are the spherical Bessel functions of the first kind Kerviche et al. (2017); Rehacek et al. (2017). The correlation functions for this basis are
$\Gamma_{q}(x_{s})=\sqrt{\sigma}\phi_{q}(x_{s})$ (11)
It was in fact shown in Sajjad et al. (2024) that this equation will exactly hold for any modal basis if we have a hard aperture in 1 dimension. For a two-dimensional hard aperture, $\sqrt{\sigma}$ will be replaced by a slightly different constant that is proportional to the square root of the area of the hard apertures. We should however point out that the results and calculations given in this paper are mostly independent of the choice of the modal basis.
III.2 The state for the two telescope set up
We now consider two apertures at positions $y_{A}=-\beta$ and $y_{B}=\beta$ for which we will also use the notation $y_{\alpha}$ with $\alpha\in\{A,B\}$. we will mostly follow the framework described in Sajjad et al. (2024), which we now summarize, but with slightly modified notation and conventions.
When we have two apertures, the aperture function $\widetilde{\psi}(y)$ gets replaced by the normalized sum of two aperture functions centered at the two telescope locations $(\widetilde{\psi}(y-y_{A})+\widetilde{\psi}(y-y_{\alpha}))/\sqrt{2}$. Now, imagine a large hypothetical imaging screen behind the two apertures spanning the entire baseline so that the light entering the two hard apertures forms an image on it. The PSF for this is the Fourier transform of this combined aperture function, just as the PSF for a single aperture is obtained by Fourier transforming the single-aperture function. Note that a shift in a function translates into a phase in the Fourier transform. Specifically, the Fourier transform of the shifted single-aperture function $\widetilde{\psi}(y-y_{\alpha})$ is $e^{-iy_{\alpha}x}\psi(x)$, where $\psi(x)$ is the single-aperture PSF obtained by Fourier transforming the unshifted function $\widetilde{\psi}(y)$. Applying this to the contributions from both apertures, the combined aperture function is
@@ III.2 The state for the two telescope set up
that is, the PSF for a single unshifted aperture times a cosine term containing information about the two-aperture system (with $\sqrt{2}$ arising from the normalization. For the special case of two hard, rect shaped apertures with the aperture function (9) centered at positions ±$\pm\beta$ in the aperture plane, we obtain the combined PSF
@@
We will refer back to this in the last section when we evaluate the performance of the measurement carried out by our protocol.
The quantum state for a single photon originating from a point source at position $x_{s}$ and entering our two-aperture system is a linear combination of starlight arriving at site $A$ or $B$
@@
$=\frac{1}{\sqrt{2}}\left(\left|\psi_{A}^{(s)}\right\rangle+\left|\psi_{B}^{(s)}\right\rangle\right)$ (14)
@@
$\left|\psi_{\alpha}^{(s)}\right\rangle=\int\text{d}xe^{-iy_{\alpha}(x-x_{s})}\psi(x-x_{s})|x\rangle,\quad\alpha\in\{A,B\}$ (15)
@@
for a single photon in mode $q$ of the telescope at site $A$, and
@@
for a single photon in mode $q$ at site $B$. Here $|0_{\alpha}\rangle\equiv\bigotimes_{i=0}^{\infty}|0_{\alpha_{i}}\rangle$ is the vacuum in all the spatial modes at site $\alpha$, in terms of which our position ket on an imaginary imaging screen spanning the entire baseline is $|x\rangle=a^{\dagger}(x)|0_{A},0_{B}\rangle$, and $a_{\alpha_{q}}^{\dagger}=\int\text{d}xe^{-iy_{\alpha}x}\phi_{q}(x)a^{\dagger}(x)$ is the creation operator for a photon in the $q$-th mode at site $\alpha$.
We will refer to $|1_{A_{q}},0_{B}\rangle$ and $|0_{A},1_{B_{q}}\rangle$ as the local modes of the telescopes at sites $A$ and $B$, respectively. The local modes of one aperture will be orthogonal to those of another one, and this has been shown in detail in Sajjad et al. (2024). Overall, the local modes of all our apertures will provide a complete orthonormal basis for our multiple aperture system.
@@
$=$ $\frac{1}{\sqrt{2}}\sum_{q=0}^{\infty}\Gamma_{q}(x_{s})\left(e^{-i\beta x_{s}}|1_{A_{q}},0_{B}\rangle+e^{i\beta x_{s}}|0_{A},1_{B_{q}}\rangle\right)$ (19)
where in the second line, we have plugged in $y_{A}=-\beta$ and $y_{B}=\beta$ for the telescope positions.
In a realistic physical set up, we will only be able to collect a finite number of spatial modes at each site. Assuming that we are collecting only the first $K$ modes, this results in the above state being projected on to the subspace limited to $10\leq q\leq K-$ 1
@@
Our goal is to use shared entanglement to measure the incoming photon in the pairwise basis introduced in Sajjad et al. (2024)
$\left|\phi_{AB,q}^{\pm}\right\rangle=\frac{1}{\sqrt{2}}\left(|1_{A_{q}},0_{B}\rangle\pm|0_{A},1_{B_{q}}\rangle\right)$ (23)
The projection of (20) onto these yield the respective probabilities
@@
$p_{q}^{-}=\eta^{2}_{q}(x_{s})\sin^{2}(\beta x_{s}),$ (25)
which are the probabilities we want to replicate with the entanglement-based protocol presented here. In the limit where we collect a large number of spatial modes i.e. $K\to\infty$, these approach $\Gamma^{2}_{q}(x_{s})\cos^{2}(\beta x_{s})$ and $\Gamma^{2}_{q}(x_{s})\sin^{2}(\beta x_{s})$, respectively, which are the probabilities obtained in Sajjad et al. (2024).
Having spelled out the ket for a photon originating from a given point source, we can write down the density operator describing the weak incoming signal, arriving at sites $A$ and $B$
@@
Where $\rho_{0,AB}=|0_{A},0_{B}\rangle\langle 0_{A},0_{B}|$ is the vacuum state across both locations which we defined above, and the one-photon state for the two aperture system is a mixture of the one-photon states associated with the two point sources
@@
where $|\psi^{(s)}_{AB,K}\rangle$ was given in (20).
III.3 The state over multiple time bins
During the measurement time, the state that describes the incoming light in a given time interval $T=tM$ is $\rho^{\otimes M}_{AB}$, where $M$ is the total number of temporal modes or time bins indexed with $m\in\{1,\ldots,M\}$. By computing this tensor product (see Appendix A for a detailed derivation), and keeping only linear terms in $\epsilon$ (in order to guarantee the presence of at most one photon over the total measuring time $T$), we obtain

### v2

@@ II Problem set up and main results
where, $|1_{A_{q}},0_{B}\rangle$ (and $|0_{A},1_{B_{q}}\rangle$, respectively) is the state of a single photon in spatial mode $q$ at site $A$ (and $B$, respectively). In Ref. Sajjad et al. (2024), it was shown how to perform this projective measurement by mixing the aforesaid two modes, for each $q$, in a 50:50 beamsplitter, followed by photon detection on both output modes, and depending upon which detector clicks, the two-mode state would be projected on to one of two states $\left|\phi_{AB,q}^{+}\right\rangle$ or $\left|\phi_{AB,q}^{-}\right\rangle$.
The next section of this paper will show in explicit terms: (1) how the above QFI-attaining projective measurement can be realized without bringing light contained in the two spatial modes $A_{q}$ and $B_{q}$ physically to mix on a beamsplitter, but by employing pre-shared entanglement stored in registers of atomic qubit memories located at each telescope site, (2) how to incorporate the compressive photon-to-memory state transfer from Ref. Khabiboulline et al. (2019a)—to load the quantum state of $1$ photon in $M$ temporal modes, into $\log_{2}(M)$ atomic qubits—while still being able to realize the underlying QFI-optimal projection (2) from Ref. Sajjad et al. (2024) on the quantum state of the incoming photon field, and (3) how to generalize this for a nonlocal entanglement-mediated realization of the QFI-optimal measurement for any quantitative imaging problem for a multi-telescope long-baseline system.
Figure 2: Schematic showing light originating at an object plane, propagating until reaching the aperture plane, with $n=2$ hard apertures each of size $\delta$, and finally propagation to an ‘image plane’, where the light could be either detected with a pixelated photon detector array, or be fed into a SPADE followed by photon detection on the sorted modes  [^2] For a long baseline system, where $\beta\gg\delta$, the image plane shown (and the ‘image-plane’ fields and quantum states considered in the paper) is a hypothetical image plane that is a Fourier transform away from the (hypothetical) compound aperture plane comprising the two hard apertures separated by a distance, and hence ‘sees’ the compound aperture’s PSF, which is a centered sinc function times a cosine. In reality, each telescope employs a local (e.g., heterodyne) detection of the locally-collected optical field, or transmits the locally-collected field in a light pipe or optical fiber to a common site where light from different telescopes are brought into an interferometer..
III Quantum state of the collected light, the ‘science photons’
In this Section, we will express the joint quantum state of the ‘science photons’, i.e., the light from a stellar scene comprising a constellation of equally-bright stars, collected at both telescope sites, across all the locally-sorted spatial modes at each telescope site and across $M$ temporal modes. We will assume the weak thermal state model described earlier, with $M\approx 1/\epsilon$ where $\epsilon\ll 1$ is the mean (thermal) photon number per mode. We will do this in three steps. In the first subsection, we write the quantum state for a single temporal mode at just one telescope site spanning all spatial modes. In the second subsection, we will generalize this to the joint quantum state of the light collected at $n=2$ telescopes over all spatial modes. Finally, in the third subsection, we write down the full quantum state across both telescopes, across all spatial modes, and $M$ temporal modes.
III.1 Quantum state for a single-aperture telescope
In the weak-thermal-source model Tsang et al. (2016a), the density operator—for a constellation of equally-bright point sources located at locations $x_{s}$ (where $s$ is a discrete index over the point sources making up the constellation)—of the state of a single temporal mode (but spanning all the spatial-modal content) at one telescope, is given by:
@@ III.1 Quantum state for a single-aperture telescope
where we have made the approximation that $\epsilon$ is small enough to warrant considering terms only up to linear in $\epsilon$ and ignore $O(\epsilon^{2})$ and higher-order terms. Here, $\rho_{0}=|0\rangle\langle 0|$ denotes the vacuum state across the collected spatial-mode span, and the single-photon state $\rho_{1}$ is a mixture of pure states of the incoming light corresponding to the one-photon states of the image-plane shifted-PSF-modes corresponding to each star in the constellation; labeled $s\in\{1,2\}$ for our two-point-source problem:
@@
In the single aperture case, we can express:
@@
where $x$ is the image-plane coordinate, $|x\rangle=a^{\dagger}(x)|0\rangle$ is the state of a single photon at position $x$ (i.e., a single-photon state of a delta-function ‘pixel mode’ in the image plane), and $a^{\dagger}(x)$ the bosonic creation operator in the continuum limit which satisfies the canonical commutation relation $[a(x),a^{\dagger}(x^{\prime})]=\delta(x-x^{\prime})$ Yuen and Shapiro (1978). Lastly, $\psi(x-x_{s})$ is the (laterally-shifted) point-spread function (PSF) corresponding to a single aperture, with $x_{s}$ being the position of the $s$-th star in the constellation.
Now, consider sorting the light collected by the aperture of the telescope in a spatial mode basis described by orthonormal mode functions $\phi_{q}(x)$ with $q=0,1,2,\ldots$. Expanding the state $\left|\psi^{(s)}\right\rangle$ defined in Eq. (5) into this spatial-mode basis gives us:
@@
is the state of a single photon in the mode $\phi_{q}(x)$ (and vacuum in all the other modes in the orthogonal complement of $\phi_{q}(x)$), and the coefficients $\Gamma_{q}(x_{s})$ are the correlation functions between the PSF and the mode functions:
@@
If one detects photons on the sorted modes, if a photon arrival is detected, the probability of detecting that photon to be in the $q$th mode is then given by: $p_{q}(x_{s})=|\Gamma_{q}(x_{s})|^{2}$.
To obtain the PSF, we need to take the Fourier transform of the aperture function $\widetilde{\psi}(y)$, where $y$ is the aperture-plane coordinate. We will consider a one-dimensional hard aperture function:
@@
where $\delta$ is the size (diameter) of the aperture. The Fourier transform of this yields the PSF
$\psi(x)=\sqrt{\sigma}\,\frac{\sin(\pi x/\sigma)}{\pi x},$ (11)
@@
While we could detect the collected field in any mode basis, one natural basis set—which we will refer to as the PSF-adapted (PAD) basis—can be constructed by taking the PSF as the fundamental mode, taking derivatives of it of increasing order and normalizing the resulting functions to obtain a set of linearly-independent (but non-orthogonal) mode functions, and finally applying a Gram-Schmidt (GS) orthogonalization process to the above set to obtain the PAD-basis mode functions. For the above hard aperture PSF, the PAD basis is the Sinc-Bessel (SB) basis, for which the mode functions are given by: $\phi_{q}(x)\equiv\sqrt{(1+2q)/\sigma}j_{q}(\pi x/\sigma)$, where $j_{q}(\pi x/\sigma)$ are the spherical Bessel functions of the first kind Kerviche et al. (2017); Rehacek et al. (2017). The correlation functions for this basis are
$\Gamma_{q}(x_{s})=\sqrt{\sigma}\phi_{q}(x_{s}).$ (12)
It was in fact shown in Sajjad et al. (2024) that Eq. (12) exactly holds for any modal basis if we have the above said hard aperture PSF in 1 dimension. For a two-dimensional hard aperture, $\sqrt{\sigma}$ in Eq. (12) will be replaced by a slightly different constant that is proportional to the square root of the area of the hard apertures. We should however point out that the results and calculations given in this paper are mostly independent of the choice of the modal basis.
III.2 Quantum state for the two telescope system
We now consider two telescopes (apertures) at positions $y_{A}=-\beta$ and $y_{B}=\beta$ for which we will also use the notation $y_{\alpha}$ with $\alpha\in\{A,B\}$. We will mostly follow the framework described in Sajjad et al. (2024), which we now summarize, but with slightly modified notation and conventions.
When we have two apertures, the aperture function $\widetilde{\psi}(y)$ gets replaced by the normalized sum of two aperture functions centered at the two telescope locations $(\widetilde{\psi}(y-y_{A})+\widetilde{\psi}(y-y_{\alpha}))/\sqrt{2}$. Now, imagine a large (hypothetical) imaging screen behind the two apertures spanning the entire baseline so that the light entering the two hard apertures forms an image on it. The PSF for this imaging system is the Fourier transform of this (compound) aperture function, just as the PSF for a single aperture is obtained by Fourier transforming the single-aperture function. A lateral shift in a function translates into multiplication by a phase in the Fourier transform. Specifically, the Fourier transform of the shifted single-aperture function $\widetilde{\psi}(y-y_{\alpha})$ is $e^{-iy_{\alpha}x}\psi(x)$, where $\psi(x)$ is the single-aperture PSF obtained by Fourier transforming the unshifted function $\widetilde{\psi}(y)$. Applying this to the contributions from both apertures, the combined aperture function comes out to:
@@ III.2 Quantum state for the two telescope system
that is, the PSF for a single unshifted aperture times a cosine term whose spatial frequency contains information about the baseline (separation) of the two-aperture system, with a $\sqrt{2}$ arising from the normalization. For the special case of two hard, rect shaped apertures with the aperture function (10) centered at positions ±$\pm\beta$ in the aperture plane, we obtain the compound PSF:
@@
where the additional subscript (h) denotes ‘hard’ apertures. We will refer back to this in the last section of the paper when we evaluate the quantum-limited performance of the measurement carried out by our protocol, limited only by this compound aperture.
The quantum state of a single photon originating from a point source at object-plane position $x_{s}$ and entering our two-aperture system is an equal linear superposition of the states corresponding to the starlight arriving at site $A$ and at site $B$:
@@
$=\frac{1}{\sqrt{2}}\left(\left|\psi_{A}^{(s)}\right\rangle+\left|\psi_{B}^{(s)}\right\rangle\right),$ (15)
@@
$\left|\psi_{\alpha}^{(s)}\right\rangle=\int\text{d}xe^{-iy_{\alpha}(x-x_{s})}\psi(x-x_{s})|x\rangle.\quad\alpha\in\{A,B\}$ (16)
@@
for a single photon in the (local) spatial-mode $q$ of the telescope at site $A$, and
@@
for a single photon in the (local) spatial-mode $q$ at site $B$. Here $|0_{\alpha}\rangle\equiv\bigotimes_{i=0}^{\infty}|0_{\alpha_{i}}\rangle$ is the vacuum in all the spatial modes at site $\alpha$, in terms of which our position ket on an imaginary imaging screen spanning the entire baseline is $|x\rangle=a^{\dagger}(x)|0_{A},0_{B}\rangle$, and $a_{\alpha_{q}}^{\dagger}=\int\text{d}xe^{-iy_{\alpha}x}\phi_{q}(x)a^{\dagger}(x)$ is the creation operator for a photon in the $q$-th mode at site $\alpha$.
We will refer to $|1_{A_{q}},0_{B}\rangle$ and $|0_{A},1_{B_{q}}\rangle$ as the single-photon states of the respective $q$-th local (shifted PAD-basis) spatial mode of the telescopes at sites $A$ and $B$, respectively. The local mode function set of one aperture will be orthogonal to those of another one, which is easy to see in the $y$-plane (see Fig. 2) considering the two local PAD-basis mode sets are mutually spatially non-overlapping (given the hard apertures). Overall, the local modes of all our apertures together provides a complete orthonormal modal basis for our multiple aperture system.
@@
$=$ $\frac{1}{\sqrt{2}}\sum_{q=0}^{\infty}\Gamma_{q}(x_{s})\left(e^{-i\beta x_{s}}|1_{A_{q}},0_{B}\rangle+e^{i\beta x_{s}}|0_{A},1_{B_{q}}\rangle\right),$ (20)
where in the second line, we have plugged in $y_{A}=-\beta$ and $y_{B}=\beta$ for the respective telescope positions.
In a realistic physical set up, we will only be able to collect a finite number of spatial modes at each site. Assuming that we are collecting only the first $K$ local PAD-basis modes, this results in the above state being projected on to the subspace limited to $10\leq q\leq K-$ 1,
@@
Our goal is to use shared entanglement to measure each temporal mode of the incoming photon field in the pairwise bases Sajjad et al. (2024):
$\left|\phi_{AB,q}^{\pm}\right\rangle=\frac{1}{\sqrt{2}}\left(|1_{A_{q}},0_{B}\rangle\pm|0_{A},1_{B_{q}}\rangle\right),$ (24)
while ensuring that the measurement strategy heralds the temporal modes that contain a photon (like the original physical strategy from Sajjad et al. (2024) of pairwise beamsplitters followed by photon detection would). The projection of (21) onto this measurement basis yields probabilities:
@@
$p_{q}^{-}=\eta^{2}_{q}(x_{s})\sin^{2}(\beta x_{s})$ (26)
for the two outcomes, which are the probabilities we want to replicate with the entanglement-based protocol to be presented in this paper. In the limit where we collect a large number of spatial modes i.e. $K\to\infty$, these approach $\Gamma^{2}_{q}(x_{s})\cos^{2}(\beta x_{s})$ and $\Gamma^{2}_{q}(x_{s})\sin^{2}(\beta x_{s})$, respectively, which are the probabilities obtained in Sajjad et al. (2024).
Having spelled out the ket for a photon originating from a given point source, we can write down the joint density operator describing the weak incoming signal, arriving at sites $A$ and $B$,
@@
where $\rho_{0,AB}=|0_{A},0_{B}\rangle\langle 0_{A},0_{B}|$ is the vacuum state across both locations across their entire local spatial-mode spans, which we defined above, and the one-photon state $\rho_{1,AB}$ for the two aperture system is a mixture of the one-photon states associated with the two point sources:
@@
where $|\psi^{(s)}_{AB,K}\rangle$ was given in (21), and the factor of $\frac{1}{2}$ arises from equal brightness of the two point sources.
III.3 Quantum state of the collected light for two telescopes over $M$ temporal modes
As described in Section II, we will be interested not only in a nonlocal realization of the projection described in Eq.(24) on each temporal mode (while heralding the presence of a photon in each temporal mode), but we also wish to do that in a way so as to allow for the compressive loading of $1$ photon in $M$ temporal modes into $\log_{2}(M)$ qubit memories per spatial mode at each site (to minimize the shared-entanglement resource consumption). We consider a measurement time interval $T$ (seconds) such that the number of temporal modes $M=WT\approx 1/\epsilon$, where $\epsilon$ is the mean photon number per temporal mode (spanning both telescopes) and $W$ (Hz) is the optical bandwidth of the source (after any spectral filtering performed by the telescope). The state of the incoming light is $\rho^{\otimes M}_{AB}$, with $m\in\{1,\ldots,M\}$ indexing the temporal modes in the time interval. By computing this tensor product (see Appendix A for a detailed derivation), and keeping only terms linear in $\epsilon$ (because the choice of $T$, hence $M$, ensures at most one photon during time $T$), we obtain: